in a class of 30 students, 7 have tape recorders, 15 have calculators, and 2 have both. how many of the 30 students have neither.
Answer by ewatrrr(24785)

7 + 15 - 2 = 20, 30-20 = 10 have neither
0r one can think of it as:
(7-2) + (15-2) + 2 = 20